OpenStudy (anonymous):

Add the following Fractions 7/9+2/7

pooja195 (pooja195):

Hi Welcome to openstudy :-) any ideas on where we would start?

OpenStudy (anonymous):

Hi no i have no clue

pooja195 (pooja195):

I see alright first thing we need to do is find an LCD (least common denominator)

pooja195 (pooja195):

In this case you would multiply by the denominators just multiply them together 9*7=?

OpenStudy (anonymous):

63

pooja195 (pooja195):

Good Thats our common denominator \[\huge~\rm~\bf~\frac{ 49+18 }{ 63 }\] 49+18=?

OpenStudy (anonymous):

soo 11/63

pooja195 (pooja195):

hmm nope try again 49+18=?

OpenStudy (anonymous):

67/63

pooja195 (pooja195):

Good and that would be our final answer

OpenStudy (anonymous):

ok thanks

pooja195 (pooja195):

yw ^_^

OpenStudy (anonymous):

i have more would u mind helping?

pooja195 (pooja195):

I can help :-)

OpenStudy (anonymous):

yay thanks

OpenStudy (anonymous):

what grade are u in?

OpenStudy (anonymous):

1 4/63

pooja195 (pooja195):

that doesnt matter ;) post your question

OpenStudy (anonymous):